OpenStudy (queenechinose):

Derivative of (ax+b)^2

OpenStudy (loser66):

Where are you stuck?

OpenStudy (thephysicsman):

well, derivative with respect to x, this is a chain rule

OpenStudy (loser66):

ok, go ahead

OpenStudy (3mar):

This is the derivative if you want.

OpenStudy (queenechinose):

Why did you get a as the first derivative of the first term? I'm always confuse about that.

OpenStudy (3mar):

Because we differentiate the (.....)^n and then differentiate what is in the (........) So the differentiation of what in the (........) is (a)

OpenStudy (queenechinose):

I can't understand. I've got the same problem a while ago. Can you explain further? Like some examples? I'm sorry if I'm being pushy on this. This is the only part I can't understand and I've got loads of Practice problems with the same concept

OpenStudy (thephysicsman):

\[\frac{ d }{ dx }(ax+b)^{2} = 2(ax+b)*\frac{ du }{ dx }\] \[u = ax+b | \frac{ du }{ dx } = ? \]

OpenStudy (queenechinose):

Yes I've got that part

OpenStudy (queenechinose):

It's a.

OpenStudy (queenechinose):

thank you